OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!!!!!!!!! Find the equation of a line with the given points. (5, 4) (5, 6)

OpenStudy (anonymous):

I think I showed you this on your other question?

OpenStudy (anonymous):

y2-y1/x2-x1=y-y2/x-x2 ok?

OpenStudy (mathstudent55):

First, find the slope of the line. Since you have two points, subtract the y-coordinates and divide by the subtraction of the x-coordinates. That gives you the slope. Once you have the slope, write y = mx + b, but write the actual value of the slope in for m. Then use the equation y = mx + b (with the real value for slope in for m), and substitute the x- and y-coordinates of one point into x and y, respectively. The only unknown left will be b. Solve for b. Then write the equation once again as y = mx + b, but substitute m and b with the numbers you found.